LOG-IN
Displaying reports 2081-2100 of 3359.Go to page Start 101 102 103 104 105 106 107 108 109 End
R&D (FilterCavity)
Print this report.
YuhangZhao - 09:49, Thursday 25 July 2019 (1511)Get code to link to this report
open-loop transfer function of filter cavity now

I upload here the open-loop transfer function of the filter cavity. The situation now we are operating with.

(gain of 7, attenuation of 7)

Images attached to this report
1511_20190725042026_figure1.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 23:30, Wednesday 24 July 2019 (1510)Get code to link to this report
Improvement of phase shifter behavior by putting rubber(2)

Aritomi and Yuhang

After seeing the effect of rubber in entry 1508, we tried to put it at another position. Considering what we reported in entry 1471, we found the resonance of 500Hz/600Hz is due to different base plate we put for phase shifter. So it seems to be an overall resonance of the mirror mount.(flag pole mode) So we decided to put a piece of rubber under the whole mirror mount. We thought it will damp the resonance of 500Hz/600Hz. The rubber and the position we put for CC1 and CC2 is shown in the attached figure 1 and 2.

After put them we increase gain little by little, we couldn't find the resonance of 500Hz/600Hz. It makes a big difference. The measurement of the optomechanical transfer function is attached.

Images attached to this report
1510_20190724162825_wechatimg480.jpeg 1510_20190724162830_wechatimg481.jpeg 1510_20190725044101_cc1.png 1510_20190725044106_cc2.png
R&D (FilterCavity)
Print this report.
NaokiAritomi - 23:10, Wednesday 24 July 2019 (1509)Get code to link to this report
Squeezing with high bandwidth phase control

[Aritomi, Yuhang]

First we put rubber under phase shifter to damp 500Hz resonance. 500Hz resonance almost disappeared with rubber and finally we could lock CC1 with 4kHz bandwidth and CC2 with 1.5kHz bandwidth. CC2 is still a bit unstable and maybe integrator is necessary for CC2.

Then we measured squeezing spectrum, but the spectrum is similar with before... 

Tomorrow we'll measure free running and closed loop phase noise of CC1,2 to estimate phase noise.

Images attached to this report
1509_20190724161041_sqz20190724.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 23:10, Wednesday 24 July 2019 (1508)Get code to link to this report
Improvement of phase shifter behavior by putting rubber(1)

Yuhang and Aritomi

We tried to put a rubber at the place where we think it can be related to resonance. First, we tried to put it between mirror/PZT holder and mirror mount. See attached figure 1 and 2.

In the measurement shown in entry 1503, CC1 has already been installed with this rubber while not for CC2. In this entry, I put a comparison of OLTF for CC2 with and without the rubber.

In the attached figure 3, it is shown the measurement with and without the rubber. We can see 4kHz resonance is damped.

In the attached figure 4, I add 12dB to measurement with rubber. The reason is we were locking with different unity gain frequency. In this case, we could compare better. From this comparison, we could see that the reduction of resonance at 4kHz gives energy to the peak at 600Hz and 7kHz and makes them higher.

Considering the position we put rubber, we think 4kHz corresponds to resonance along the direction of mirror/PZT holder axis.

Images attached to this report
1508_20190724161037_wechatimg478.jpeg 1508_20190724161042_wechatimg479.jpeg 1508_20190724161207_cc22.png 1508_20190724161212_ccc.png
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 16:18, Wednesday 24 July 2019 (1506)Get code to link to this report
Preliminary work on test mass feedback to reduce cavity motion

I injected noise on lenght d.o.f of the end mirror when FC is locked and measured its TF to legth oplev and correction signal to the laser piezo.

The TFs look similar (pic 1) even if they are not very clear and the coherence is quite low. (pic 2).

The noise amplitude was 10000 count (~3V) and the driving matrix was such that only vertical coils where used. (It should couple only with pitch)

Pic 3 shows the spectra of the other d.o.fs. The damping loops were closed at that time, but pitch shows a motion larger that usual.

Images attached to this report
1506_20190724085244_zexc240719p1.png 1506_20190724085300_zcorrcohe240719.png 1506_20190724092628_zcorrspe240719.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 11:51, Wednesday 24 July 2019 (1505)Get code to link to this report
The holes need to be made on the board to cover bench/length of homodyne power line

1. For main laser power line. The board should have a hole 11.1cm away from border and 4cm high.

2. For green injection, the beam is at the edge of the board. Height of 7.55cm

3. For IR injection, the beam is 37.2 holes(each hole 2.5cm)  away from main laser side.

4. length of homodyne power cable should be 5m

(Please look at attached figure to make it clear about where should be the holes)

R&D (FilterCavity)
Print this report.
NaokiAritomi - 23:30, Tuesday 23 July 2019 (1504)Get code to link to this report
Investigation of squeezing spectrum

[Aritomi, Yuhang]

First we measured dark noise of homodyne since we had large 50Hz noise in shot noise spectrum. The dark noise already had large 50Hz noise and we found that when we removed ground floating of spectrum analyzer we are using, these peaks disappeared (Pic 1).

For CC, we checked noise in SHG, GRMC (entry 1514). Then we tried to lock CC1 and 2. For CC1, at the beginning of today we could lock CC1 in stable region between 500Hz and 7kHz resonance, but at the end of today, we couldn't find stable region somehow. For CC2, we couldn't lock it stably so we put rubber to CC2 mirror holder and 4kHz peak is damped (entry 1508), but we still cannot lock it because of 7kHz resonance. 4kHz resonance should be resonance of mirror holder, but we don't know where 7kHz resonance comes from.

Then we measured squeezing with 50mW green (Pic 2) when both CC loops are stable. Squeezing is ~5dB and anti-squeezing is ~20dB. Squeezing spectrum was not very good due to phase noise. We also measured it with 40mW green (OPO:7.169kOhm, PLL:150MHz). Squeezing spectrum is more clear while squeezing level is almost same.

Images attached to this report
1504_20190723163041_darknoise.png 1504_20190723163045_squeezing.png
Comments related to this report
YuhangZhao - 13:07, Thursday 25 July 2019 (1514)

Here I attach the measurement of SHG and GRMC transmission noise spectrum. 

The difference is I measured the noise spectrum of SHG transmission at a wrong place. (Actually, I was measuring the locking noise of SHG in the entry 1501)

However, we really have a worse GRMC transmission noise. Especially, there are peaks from the 4-10kHz region. They are responsible for the region where we have more noise in the squeezing measurement.

I will try to understand why there is a factor between the old and new measurement.

Images attached to this comment
1514_20190725060914_shgn.png 1514_20190726035743_150120190723050014mz.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 19:28, Tuesday 23 July 2019 (1503)Get code to link to this report
Measurement of both coherent control loops transfer function

Aritomi and Yuhang

We measured the open-loop transfer functions of CC1 and CC2 yesterday.

We could see from these two figures that 0dB line cross 500Hz and 7kHz for CC1. While 0dB line cross 600Hz, 4kHz and 7kHz peaks at the same time for CC2.

These peaks make locking not easy to achieve. We are thinking about the following things to try:

  • to dump resonance by putting rubber
  • to use a notch filter to not excite the resonance
  • to glue PZT and mirror on the phase shifter
Images attached to this report
1503_20190723122752_cc1.png 1503_20190723122757_cc2.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 18:31, Tuesday 23 July 2019 (1502)Get code to link to this report
Comment to PDH lock (Click here to view original report: 1495)

I tried to improve the lock stability by adjusting the gain of SR560.
However, I could not improve so much.
In the end, the lock last only 0.5sec.

So we need circuits for feedback control to achieve stable lock.
Anyway, we can lock the laser to temporal silicon cavity.

R&D (FilterCavity)
Print this report.
YuhangZhao - 12:00, Tuesday 23 July 2019 (1501)Get code to link to this report
Measurement of noise spectrum of SHG and GRMC

To confirm if we have more green phase noise(free-running), I performed the measurement of noise spectrum of SHG and GRMC transmission again.

The comparison is with entry 1276.

We could see we have much more noise now!

Images attached to this report
1501_20190723050010_shg.png 1501_20190723050014_mz.png
R&D (FilterCavity)
Print this report.
YuhangZhao - 09:52, Tuesday 23 July 2019 (1500)Get code to link to this report
Comment to Recovery of 5dB squeezing (Click here to view original report: 1499)

Add some points,

  • We observed the unstable lock of green phase. And I think it is related to the unstable of squeezing measurement.
  • From the phase control behavior, the phase noise(free-running) seems to be higher than before. Need to be confirmed.
  • The green phase lock has two main resonance 500Hz and 7kHz. In the beginning, we could find a gain region between these two oscillation and lock green phase. But at the end of yesterday, we couldn't find this region when we change the gain of control servo.
  • The error signal for the green phase lock usually change. As pointed out by Aritomi-san, once it is due to the misalignment of green(green power was changed). Actually, we observed a second change of error signal of green phase lock. We check PLL lock, green power(also GRMC alignment), OPO temperature. But they were fine. We should use more time to investigate what makes it change and try to avoid it.
R&D (FilterCavity)
Print this report.
NaokiAritomi - 23:44, Monday 22 July 2019 (1499)Get code to link to this report
Recovery of 5dB squeezing

[Aritomi, Yuhang]

First we made LO DC balanced and aligned LO and BAB into AMC. DC balance of BAB is 10mV. We measured shot noise of LO (Pic 1). Although there are large 50Hz peak and its harmonics , homodyne is shot noise limited above 20Hz. To remove the 50Hz peak, we floated ground of homodyne, but the shot noise was same. We also measured shot noise spectrum when LO and CC are injected into homodyne. Shot noise with LO and CC was same as shot noise with LO only.

Then we locked green and IR phase with SR560 (bandwidth  ~ 100Hz) and found squeezing spectrum is very bad. So we locked green and IR phase with Pierre's servo (bandwidth ~ 1kHz) and finally could recover the 5dB squeezing (Pic 2). 

We thought green power is 50mW, but actually green power was smaller than 50mW due to misalignment of GRMC at that time.

, homodyne is shot noise limited above 20Hz. To remove the 50Hz peak, we floated ground of homodyne, but shot noise was same. We also measured shot noise spectrum when LO and CC are injected into homodyne. Shot noise with LO and BAB was same as shot noise with LO only
First we made LO DC balanced and aligned LO and BAB into AMC. DC balance of BAB is 10mV.
We measured shot noise of LO (Pic 1). Although there are large 50Hz peak and its harmonics
, homodyne is shot noise limited above 20Hz. To remove the 50Hz peak, we floated ground of homodyne, but shot noise was same. We also measured shot noise spectrum when LO and CC are injected into homodyne. Shot noise with LO and BAB was same as shot noise with LO only.
Images attached to this report
1499_20190722164352_img7694.jpeg 1499_20190722164359_201907022.png
Comments related to this report
YuhangZhao - 09:52, Tuesday 23 July 2019 (1500)

Add some points,

  • We observed the unstable lock of green phase. And I think it is related to the unstable of squeezing measurement.
  • From the phase control behavior, the phase noise(free-running) seems to be higher than before. Need to be confirmed.
  • The green phase lock has two main resonance 500Hz and 7kHz. In the beginning, we could find a gain region between these two oscillation and lock green phase. But at the end of yesterday, we couldn't find this region when we change the gain of control servo.
  • The error signal for the green phase lock usually change. As pointed out by Aritomi-san, once it is due to the misalignment of green(green power was changed). Actually, we observed a second change of error signal of green phase lock. We check PLL lock, green power(also GRMC alignment), OPO temperature. But they were fine. We should use more time to investigate what makes it change and try to avoid it.
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 19:33, Monday 22 July 2019 (1497)Get code to link to this report
Comment to PDH lock (Click here to view original report: 1495)

I reduced the gain to 2*10^2, and tried to lock.
The oscilation was improved than higher gain one, but the lock last only 0.1sec and DC REFL dip was ~0.1mW.

Anyway, it seems that it is possible to lock TEM00 mode.
Actually, feeding back to laser PZT induces laser power fluctuation.
Therfore, ISS is needed in this experiment.

Images attached to this comment
1497_20190722122128_20190722lock2.jpg
KAGRA MIR (Polarization)
Print this report.
SimonZeidler - 11:50, Monday 22 July 2019 (1496)Get code to link to this report
Polarization measurements on GT-ETMX started

Matteo, Simon

On Friday, we started to upgrade the absorption map to be capable of taking maps on the homogeneity of mirror-substrates in terms of their polarization.
The basic idea is to use a PBS nad two photodiodes (PSDs in our case) to distinguish between S and P polarization. Actually, we recognized that the incoming beam is 90% P and 10% S polarized so that we decided to put another PBS also in the incoming beam in addition to a half-wave plate which we use to check the differences in S and P polarization and the calibration of the system.

After some first checks, we found out that the PSD designated for sensing the S-polarization is very sensible to any scattered light which is automatically an issue due to the many ND filters we have to use for not saturating the AC channel (which is fed by the S-pol PSD). Therefore, we took long time to cover the path from the PBS to the PSD. We finally took a beam-pipe which does the job quite well (see attached pictures).

Over the weekend, we took two first maps. One with incoming-beam fully P-polarized and the other one having the half-wave plate rotated by 30 degrees. Qualitative pictures can already be calculated but we still need to calibrate the DC signal properly to calculate also a quantitative number.

Images attached to this report
1496_20190722044926_44.jpg 1496_20190722044942_50.jpg 1496_20190722045002_map1907201973344pm.png 1496_20190722045005_map2007201970434pm.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 09:14, Monday 22 July 2019 (1495)Get code to link to this report
PDH lock

This morning, I again tried to lock TEM00 mode.
The attached picture shows when the laser seemed locked (the label is same as before).
It seemes something is fed back to the laser, and it last several tens of seconds.
Actually, it is oscillating, and the dip is so small, so it needs to improve the alignment, gain, and some other things for more stable lock.

Images attached to this report
1495_20190722014848_20190722lock.jpg
Comments related to this report
SatoshiTanioka - 18:31, Tuesday 23 July 2019 (1502)

I tried to improve the lock stability by adjusting the gain of SR560.
However, I could not improve so much.
In the end, the lock last only 0.5sec.

So we need circuits for feedback control to achieve stable lock.
Anyway, we can lock the laser to temporal silicon cavity.

SatoshiTanioka - 19:33, Monday 22 July 2019 (1497)

I reduced the gain to 2*10^2, and tried to lock.
The oscilation was improved than higher gain one, but the lock last only 0.1sec and DC REFL dip was ~0.1mW.

Anyway, it seems that it is possible to lock TEM00 mode.
Actually, feeding back to laser PZT induces laser power fluctuation.
Therfore, ISS is needed in this experiment.

Images attached to this comment
1497_20190722122128_20190722lock2.jpg
R&D (FilterCavity)
Print this report.
NaokiAritomi - 07:14, Monday 22 July 2019 (1494)Get code to link to this report
DC balance of LO and BAB

[Aritomi, Yuhang]

This is work on July 20th. First we made LO and BAB overlap and more or less DC balanced at homodyne. CMRR at 1 kHz is 82 dB which is similar to before. Then we measured shot noise and squeezing spectrum (data is not saved). Shot noise is worse below 100Hz and squeezing spectrum was very bad. 

After that we found large 7 MHz peak at homodyne DC signal (attached pictures). LO and CC before homodyne BS is 1.2 mW and 8 uW respectively. This 7 MHz seems beat between LO and CC.

Images attached to this report
1494_20190722001354_7mhzoscilloscope.jpg 1494_20190722001400_7mhzspectrum.jpg
R&D (FilterCavity)
Print this report.
NaokiAritomi - 06:38, Monday 22 July 2019 (1493)Get code to link to this report
Modification of optical layout for BAB flipping mirror

[Aritomi, Yuhang]

This is work on July 19th. In current setup, we only have one steering mirror and one lens for alignment of BAB going to homodyne and it's difficult to align BAB into AMC. So we changed optical layout so that we can have three steering mirrors for alignment of BAB. I updated optical layout of our bench. After this modification, we could align BAB to AMC easily. Re-alignment of IR to FC after this modification is ongoing.

Images attached to this report
1493_20190721233739_41.png 1493_20190721234405_scheme43.png
R&D (Cryogenic)
Print this report.
SatoshiTanioka - 18:57, Sunday 21 July 2019 (1492)Get code to link to this report
TEM00 lock trial

Today I tried to lock TEM00 mode with 10^3 gain.
The attached pictures show DC REFL (blue), error signal (green), and feedback signal (yellow).

Actually, the lock was not stable; it last only ~10msec.
This seemes due to the fluctuation of feedback siganl which oscillated about 20Hz.
Also the error signal oscillates about this frequency, and this hinders me to lock.
I will investigate from where this oscillation comes.

Images attached to this report
1492_20190721114307_201907212.jpg 1492_20190721114315_201907211.jpg 1492_20190721114319_201907213.jpg
R&D (FilterCavity)
Print this report.
EleonoraCapocasa - 09:47, Saturday 20 July 2019 (1490)Get code to link to this report
Comment to The measurement of PLL phase noise for CC and PP when filter cavity is locked or not locked (Click here to view original report: 1482)

The excess of high frequency noise measured last summer when the cavity is locked (see entry #903) seems disapeared. This is good but the reason is not clear to me.

Note that in order to sum the line into the PZT  for the calobration we have reconnected the ramp potentiometer which was disconnected last summer,  following Pierre's advice,  in order to reduce the rampeato noise (see entry #875).

We have left it connected for now. It would be interesting to compare the difference in the error signal spectrum when it is connect and disconnected as done last summer (entry #883).

R&D (FilterCavity)
Print this report.
NaokiAritomi - 08:12, Saturday 20 July 2019 (1489)Get code to link to this report
Comment to The measurement of filter cavity locking noise (Click here to view original report: 1486)

Can you do the same measurement with gain other than 10 (when error signal noise is larger)?